# Machine learning for structure-guided materials and process design

### v1

@@ 4.1 Metal forming process simulation
The metal forming process simulation [9] used in this study applies a deformation $\hat{\boldsymbol{F}}$
@@
with $\boldsymbol{R}$ being a rotation matrix that describes one out of 25 possible loading directions. The deformation $\widetilde{\boldsymbol{F}}$ is defined using orthogonal basis vectors $\boldsymbol{e}_{i}$ and the operator $\otimes$ for the outer product
@@
with the slip rates $\dot{\gamma}^{(\alpha)}$ on slip system $\alpha$ that is defined by the slip plane normal $\boldsymbol{n}^{(\alpha)}$ and the slip direction $\boldsymbol{m}^{(\alpha)}$. The slip rates are calculated by a phenomenological power law. The crystal reorientation is calculated by applying a rigid body rotation derived from the polar decomposition of $\boldsymbol{F}_{\mathrm{e}}$ to the original orientation [27, 11].
4.2 Texture representation and distance measure
@@
Texture is represented by using the histogram-based description introduced by Dornheim et al. [9]. The orientation histogram (soft-assignment factor of three) consists of 512 orientation bins that are nearly uniformly distributed in the cubic-triclinic fundamental zone, which were created using the software neper [28, 29]. In contrast to the original MEG-SGGPO approach, crystallographic texture distances are measured using the Sinkhorn distance applied to the histogramm representation [11]. Specifically, the Sinkhorn distance is an efficient implementation of the Earth Movers distance that measures the least amount of work necessary to transform one histogram into the other [30, 31]. Therefore, local orientation distances [32] encoded in the histograms are taken into account, in contrast to the originally proposed Chi-Squared distance, which is basically a bin-wise comparison of two histograms.
4.3 The SMTLO approach
The SMTLO approach described in [10] aims to identify diverse microstructures for given desired properties (defined by a target region). It consists of a multi-task learning model, which is grounded on an encoder $f_{\mathrm{x\text{-}enc}}$ that transforms a microstructure representation $\boldsymbol{x}$ into a lower dimensional latent feature space representation $\boldsymbol{z}$
$\boldsymbol{z}=f_{\mathrm{x\text{-}enc}}(\boldsymbol{x},\boldsymbol{\theta}_{\mathrm{x\text{-}enc}}),$ (8)
with the trainable parameters $\boldsymbol{\theta}_{\mathrm{x\text{-}enc}}$. The encoder is equipped with three heads, each of which is designed to solve a specific learning task. These are
  1. 1.
    $\boldsymbol{x}^{\prime}=f_{\mathrm{x\text{-}dec}}(\boldsymbol{z},\boldsymbol{\theta}_{\mathrm{x\text{-}dec}}),$ (9)
    with the trainable parameters $\boldsymbol{\theta}_{\mathrm{x\text{-}dec}}$. The encoder-decoder part is trained using a loss term that minimizes the Sinkhorn distance between the original microstructure $\boldsymbol{x}$ and its reconstruction $\boldsymbol{x}^{\prime}$ [11]
    $\mathscr{L}_{1}=\mathscr{D}_{\mathrm{sh}}(\boldsymbol{x},\boldsymbol{x}^{\prime}).$ (10)
  2. 2.
    a prediction head $f_{\mathrm{p\text{-}pred}}$ to infer material properties $\hat{\boldsymbol{p}}$
    $\hat{\boldsymbol{p}}=f_{\mathrm{p\text{-}pred}}(\boldsymbol{z},\boldsymbol{\theta}_{\mathrm{p\text{-}pred}}),$ (11)
    with the trainable parameters $\boldsymbol{\theta}_{\mathrm{p\text{-}pred}}$. The sample-wise loss term is defined by the mean squared error between predicted and true properties
    $\mathscr{L}_{2}=\frac{1}{n_{\mathrm{p}}}\sum_{i}^{n_{\mathrm{p}}}(p_{i}-\hat{p}_{i})^{2},$ (12)
    with the number of properties $n_{\mathrm{p}}$.
  3. 3.
    $\boldsymbol{z}^{\prime}=f_{\mathrm{z\text{-}dec}}(f_{\mathrm{z\text{-}enc}}(\boldsymbol{z},\boldsymbol{\theta}_{\mathrm{z\text{-}enc}}),\boldsymbol{\theta}_{\mathrm{z\text{-}dec}}),$ (13)
    with the encoder and decoder parts $f_{\mathrm{z\text{-}enc}}$ and $f_{\mathrm{z\text{-}dec}}$ and the trainable parameters $\boldsymbol{\theta}_{\mathrm{z\text{-}enc}}$ and $\boldsymbol{\theta}_{\mathrm{z\text{-}dec}}$, respectively. The auto-encoder is trained using a mean squared error loss function
    $\mathscr{L}_{3}=\frac{1}{n_{\mathrm{z}}}\sum_{i}^{n_{\mathrm{z}}}(z_{i}-z^{\prime}_{i})^{2},$ (14)
    with $n_{\mathrm{z}}$ dimensions of the latent feature space.
A neural network is used to solve the above mentioned learning tasks, which is trained using a combined loss function
$\mathscr{L}_{\mathrm{MTL}}=W_{1}\mathscr{L}_{1}+W_{2}\mathscr{L}_{2}+W_{3}\mathscr{L}_{3},$ (15)
with individually weighted loss terms using the parameters $W_{1}=0.2$, $W_{2}=0.06$, and $W_{3}=0.04$.
To ensure that the original microstructure distance is preserved in the latent feature space, a Siamese training method is applied [33]. This involves training two twin models simultaneously, with shared weights, and adding a further loss term that applies to the latent feature space. This preservation loss writes
$\mathscr{L}_{4}=(\mathscr{D}_{\mathrm{sh}}(\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)})-\mathrm{dist}(\boldsymbol{z}^{(1)},\boldsymbol{z}^{(2)}))^{2},$ (16)
with the absolute distance in the latent feature space
$\mathrm{dist}(z_{i}^{(1)},z_{i}^{(2)})=\frac{1}{n_{\mathrm{z}}}\sum_{i}^{n_{\mathrm{z}}}|z_{i}^{(1)}-z_{i}^{(2)}|.$ (17)
The superscripts (1) and (2) are used to distinguish between the two simultaneously trained twins parts. The overall loss function writes
$\mathscr{L}=\mathscr{L}_{\mathrm{MTL}}+W_{4}\mathscr{L}_{4}+\lambda\Omega(\boldsymbol{\theta}),$ (18)
with the regularization term $\Omega(\boldsymbol{\theta})$, $\lambda=10^{-6}$, and $W_{4}=0.7$.
In order to identify microstructures with desired properties, we use a genetic optimizer [34] to search for candidate microstructures in the latent feature space generated by the Siamese multi-task learning model. The optimizer minimizes three cost terms: $\mathscr{C}_{1},\mathscr{C}_{2},\mathscr{C}_{3}$ [10]. $\mathscr{C}_{1}$ minimizes the distance of the predicted property $\hat{\boldsymbol{p}}$ for a candidate microstructure $\boldsymbol{z}^{*}$ to the target region. $\mathscr{C}_{2}$ penalizes candidate microstructures that are not inside the region of known microstructures. $\mathscr{C}_{3}$ enforces candidate microstructures being as far away as possible from each other in the actual population. The overall objective function writes
$\mathscr{C}=V_{1}\mathscr{C}_{1}+V_{2}\mathscr{C}_{2}+V_{3}(1+\mathscr{C}_{3}),$ (19)
with individual weights $V_{1}=0.75$, $V_{2}=0.05$, and $V_{3}=0.2$.
4.4 The MEG-SGGPO approach
In reinforcement learning, an agent learns to take optimal decisions by interacting with its environment in a sequence of discrete time steps $t$. The MEG-SGGPO approach [9] is tailored for efficient learning in the case of multiple equivalent target microstructures. Given a set $\mathcal{G}$ of near-optimal microstructures $s^{(\text{g})}\in\mathcal{G}$, an initial microstructure $s_{0}$, and a microstructure generating process $p(s_{t},a_{t})=s_{t+1}$, the goal of the approach is to find a sequence of processing parameters $a_{t}\in A$ that guide the process towards manufacturing the best reachable microstructure $s^{*}\in\mathcal{G}$.
MEG-SGGPO is based on deep Q-networks[35] and its extensions [36, 37, 38]. Deep Q-networks are derived from Q-learning, where the so-called Q-function is approximated from data gathered by the reinforcement learning agent while following a policy $a=\pi(s)$. The Q-function models the expected sum of future reward signals for state action pairs $(s_{t},a_{t})$. In MEG-SGGPO the Q-function is generalized to also take the goal microstructure into account and is, in its recursive form, defined as
$\mathcal{Q}(s_{t},a_{t},s^{(\text{g})})=\mathbb{E}_{P,\pi}\Big{[}R(s_{t},s_{t+1},s^{(\text{g})})+\max_{a_{t+1}\in A}\mathcal{Q}(s_{t+1},a_{t+1},s^{(\text{g})})\Big{]},$ (20)
where $R(s_{t},s_{t+1},s^{(\text{g})})$ is a per goal microstructure pseudo reward function
$R(s_{t},s_{t+1},s^{(\text{g})})=\dfrac{1}{\mathscr{D}(s_{t+1},s^{(\text{g})})}-\dfrac{1}{\mathscr{D}(s_{t},s^{(\text{g})})},$ (21)
which is based on a microstructure distance function $\mathscr{D}$.
The process design task is a two-fold optimization problem to(i) identify the best reachable goal microstructure $s^{*}$ and (ii) learn the optimal policy $\pi^{*}$. MEG-SGGPO addresses this by using the generalized functions defined above in a nested reinforcement learning procedure. At the beginning of each episode (i.e. an execution of the process during learning), the agent uses the generalized Q-function to identify the targeted microstructure $s^{\prime}$ from $\mathcal{G}$ and determines the processing parameters during the episode. In a first step, an estimation of the best reachable goal microstructure $\bar{s^{*}}$ is extracted from the current $\mathcal{Q}$ estimation by
$\bar{s^{*}}=\operatorname*{arg\,max}_{s^{(\text{g})}\in\mathcal{G}}\Big{[}\mathcal{V}(s_{0},s^{(\text{g})})+\dfrac{1}{\mathscr{D}(s_{0},s^{(\text{g})})}\Big{]},$ (22)
where the generalized state-value function $\mathcal{V}$ is defined as
$\mathcal{V}(s,s^{(\text{g})})=\max_{a\in A}\mathcal{Q}(s,a,s^{(\text{g})}).$ (23)
The targeted microstructure $s^{\prime}$ is chosen in an $\epsilon$-greedy approach [39], where the agent chooses random targets and processing parameters in an $\epsilon$ fraction of the cases ($0\leq\epsilon\leq 1$) and optimizes targets and processing parameters in the remaining cases. The optimal processing path is identified simultaneously in an inner loop by using the Q-learning approach.

### v2

@@ 2.2.2 Metal forming process simulation
In this study, we use the metal forming process simulation as is described in [10]. The simulation applies a deformation $\hat{\boldsymbol{F}}$
@@
with $\boldsymbol{R}$ being a rotation matrix that describes one out of 25 possible loading directions. The deformation $\widetilde{\boldsymbol{F}}$ is defined using orthogonal basis vectors $\boldsymbol{e}_{i}$ and the operator $\otimes$ for the dyadic product
@@
with the slip rates $\dot{\gamma}^{(\alpha)}$ on slip system $\alpha$ that is defined by the slip plane normal $\boldsymbol{n}^{(\alpha)}$ and the slip direction $\boldsymbol{m}^{(\alpha)}$. The slip rates are calculated by a phenomenological power law. The crystal reorientation is calculated by applying a rigid body rotation derived from the polar decomposition of $\boldsymbol{F}_{\mathrm{e}}$ to the original orientation, see [37, 12]. For the metal forming process simulation, the same material model parameters have been used as in [10].
For the purpose of this study, we use the above described material model to evaluate the Young’s modulus $E$ and an anisotropic property $\widetilde{R}$ in three orthogonal directions after a process run. The Young’s modulus is calculated using the slope of the stress-strain curve in the elastic regime, that results after applying uniaxial tension. $\widetilde{R}$ is inspired by the Lankford coefficients in sheet metal forming and is calculated as the ratio between transverse strain and the strain in the loading direction in the plastic regime, also after applying uniaxial loading.
2.2.3 Crystallographic texture representation and texture distance measure
@@
Texture is represented by using the histogram-based description introduced by Dornheim et al. [10]. The orientation histogram (here, we used a soft-assignment factor of 3) consists of 512 orientation bins that are nearly uniformly distributed in the cubic-triclinic fundamental zone. The bin orientations were created using the software neper [38, 39]. In contrast to the original MEG-SGGPO and SMTLO approach, in this work, crystallographic texture distances are measured using the Sinkhorn distance that is applied to the histogramm representations [12]. Specifically, the Sinkhorn distance is an efficient implementation of the Earth Movers distance that measures the least amount of work necessary to transform one histogram into the other [40, 41]. Therefore, local orientation distances [42] encoded in the histograms are taken into account, in contrast to the originally proposed Chi-Squared distance, which is basically a bin-wise comparison of two histograms.
3 Results
3.1 Materials and process design task
In the following sections, it is shown how near-optimal crystallographic textures are identified for given elastic and anisotropy properties using the SMTLO approach. The properties are in particular the Young’s moduli $E_{i}$ and anisotropy measures $\widetilde{R}_{i}$, both in three orthogonal directions.
The basis for applying the SMTLO approach is a data set of 76980 samples, composed of textures and corresponding properties. 2D projections of the training data distribution, generated by the metal forming process simulation, are shown in Figure 4 including the region delineating the desired properties (target region). In this study, the target region is centered at $E_{11},E_{22},E_{33}=214,214,221$ GPa and $\widetilde{R}_{23},\widetilde{R}_{12},\widetilde{R}_{13}=0.65,0.685,0.885$. Its width equals $2$ GPa for $E_{i}$ and $0.2$ for $\widetilde{R}_{i}$. Both $E_{i}$ and $\widetilde{R}_{i}$ are determined in three orthogonal directions at the material point at the end of the process. $E_{i}$ is given by the slope in the elastic regime, while $\widetilde{R}_{i}$ is calculated by the ratio between transverse and tensile strain in the plastic regime when applying uniaxial tension. In Figure 4, we highlighted training data points that are already located inside the target region. We use this set of $80$ textures as benchmark set.
3.2 Materials design using SMTLO
The SMTL model is realized via feedforward neural networks with $tanh$ activation functions and are implemented based on the Pytorch API [43]. For hyperparameter optimization the random search method [44] is utilized using 5-fold cross-validation. The following hyperparameters have been chosen: The Glorot Normal method [45] is used for weight initialization and the Adam optimizer [46] is used with the following parameters: learning rate = 0.001, weight decay = $10^{-6}$, batch size = 128. The SMTL model is trained for 200 epochs, where the best intermediate result of the test set is retained to prevent over fitting and to apply early stopping [47]. Before the model is trained, the loss terms are scaled to values between 0 and 1 in order to make them comparable. The following weights for the scaled loss terms in Eq. (11) were based on hyper parameter optimization: $\mathscr{W}_{\mathrm{recon}}=0.2$, $\mathscr{W}_{\mathrm{regr}}=0.06$, and $\mathscr{W}_{\mathrm{valid}}=0.04$, and $\mathscr{W}_{\mathrm{pres}}=0.70$. For the cost function of the optimizer (defined in Eq. (12)) that we need after having a trained siamese multi-task learning model, we use the weights $\mathscr{V}_{\mathrm{prop}}=0.75$, $\mathscr{V}_{\mathrm{valid}}=0.05$, and $\mathscr{V}_{\mathrm{divers}}=0.2$. The results for the properties prediction are given by the the mean absolute error (MAE) between the true and predicted Young’s moduli and $\widetilde{R}$-values: $\mathrm{MAE}_{\mathrm{E}}$ = 0.368 [GPa] and $\mathrm{MAE}_{\mathrm{\widetilde{R}}}$ = 0.032 [-]. The result of the distance preservation is measured by the coefficient of determination $R^{2}$, between the Sinkhorn distance of two input textures and the $l_{1}$ distance of their corresponding latent feature vectors: $R^{2}(\mathscr{D}_{\mathrm{sh}}(\boldsymbol{x_{L}},\boldsymbol{x_{R}}),l_{1}(\boldsymbol{z}_{L},\boldsymbol{z}_{R}))=90.22[\%].$
Using the SMTLO approach (with the trained multi-task learning model as described above), we were able to identify a set of 175 near-optimal textures. As we aim to identify a preferably diverse set of textures, we depict their pairwise distances in Figure 5 and compare them to the benchmark set. As one can easily see, the SMTLO approach is able to find a set of textures that differ more to each other than the benchmark set.
For the subsequent process design step, we chose ten goal textures from the set of identified textures that differ strongly from each other [48]. The distribution of the chosen goal textures in properties space is shown in Figure 6. The properties of the goal textures are mainly located inside the target region, with some exceptions that we tolerate for now and explain later in Section 4. To show the differences between the goal textures, in Figure 7, four textures are depicted exemplary as pole figure plots. While the pole figure intensities are similar for all textures, the shape of the represented orientation distribution differs strongly.
Figure 6: Projections of the properties space showing the target region and the calculated properties of the ten chosen goal textures $\check{g}^{0}$ to $\check{g}^{9}$
3.3 Process design using MEG-SGGPO
Experiments are conducted with the following hyperparameters: Deep Q-learning as described in Section 2.1.3 as basic algorithm where the target-network is updated every $n_{\theta}=50$ time-steps. Q-networks with hidden layer sizes of [128, 256, 256, 128], layer normalization and ReLU activation functions. The learning process starts after 100 control-steps. The networks are trained after each control-step with a mini-batch of size 32. The Adam optimizer [cite] is used for neural network training, with a learning rate of $5e^{-4}$. An $\epsilon$-greedy policy, with an initial exploration rate $\epsilon_{0}=0.5$ and the final exploration-rate $\epsilon_{f}=0.0$, with $n_{\epsilon}=390$.
Once a diverse set of crystallographic textures has been identified, MEG-SGGPO is used to guide the metal forming process to the best reachable texture. For solving the identification problem, the MEG-SGGPO approach is allowed to conduct $400$ process runs, so-called episodes. The evolution of the distance between the produced and the chosen goal texture is depicted over the episodes in Fig. 8. Initially, the reinforcement learning agent attempts to identify processing paths that target randomly selected goal textures. As the learning process progresses, the agent focuses increasingly on producing goal texture $\check{g}^{2}$.
For comparison, Figure 9 presents pole figure plots of the goal texture goal texture $\check{g}^{2}$ and the produced texture $\check{g}^{2^{\prime}}$ after 400 episodes. Visually, the produced texture is highly similar to the goal texture with intensity peaks at the same positions and of a similar magnitude. This also holds for areas not covered by orientations at all. To quantify the effectiveness of the approach, we compute the properties resulting from the produced crystallographic texture and compare them with the desired ones. The results, listed in Table 1, show that five of the properties ($E_{11}$, $E_{22}$, $E_{33}$, $\widetilde{R}_{12}$ and $\widetilde{R}_{13}$) are very close to the target region, while $\widetilde{R}_{23}$ lies inside.
Figure 9: Pole figure plots (001), (110), (111) of the targeted goal texture $\check{g}^{2}$ (top) and the produced texture $\check{g}^{2^{\prime}}$ (bottom)
Target property Value Distance to target region Unit
$E_{11}$ 217.5 2.5 GPa
$E_{22}$ 216.6 1.6 GPa
$E_{33}$ 222.2 0.3 GPa
$\widetilde{R}_{23}$ 0.744 0 -
$\widetilde{R}_{12}$ 0.831 0.046 -
$\widetilde{R}_{13}$ 1.058 0.073 -
In summary, by sequentially applying the SMTLO and the MEG-SGGPO approach, we were able to successfully produce a crystallographic texture with desired properties in a simulated metal forming process. The resulting properties fall within the acceptable range from an engineering point of view, albeit at the border of the target region defined in properties space.
4 Discussion
The results presented demonstrate the effectiveness of the SMTLO approach in identifying sets of near-optimal crystallographic textures that exhibit desired properties in a given target region. The set of textures identified by the approach is shown to be more diverse than the benchmark set obtained from the training data set. However, the properties of the identified textures are not completely inside the target region, as can be seen in Fig. 6.
There are two reasons for this: First, the formulation of the optimization objective in the SMTLO approach allows for a trade-off between (i) finding crystallographic textures with properties inside the target region, (ii) identifying possibly diverse textures and, (iii) guaranteeing that texture can be produced by the process under consideration. Second, due to prediction errors of the underlying machine learning model, the SMTLO approach identifies textures whose predicted properties lie inside the target region, while some of the true properties (calculated by the numerical simulation on the basis of the reconstructed textures) slightly lie outside. Both of these issues can be addressed by modifying the objective function of the SMTLO optimizer, defined by Equation 12. For instance, targeting the center of the properties region, instead of its bounds can mitigate both issues. Additionally, the second issue can be addressed by enhancing the machine learning model, for example, by increasing the amount of training data.
Taking a diverse subset of the identified textures as goal textures, the MEG-SGGPO approach guides the metal forming process closely to one of the chosen goal textures. Although the reinforcement learning agent was not able to reproduce one of the goal textures exactly (which is challenging due to the many possible processing paths), the measured distance between the targeted texture $\check{g}^{2}$ and the produced one $\check{g}^{2^{\prime}}$ is sufficiently low ($\mathscr{D}_{\mathrm{sh}}=0.031$) when relating it to the distance to the nearest neighbor in the training data set ($\mathscr{D}_{\mathrm{sh}}=0.033$) and to the furthest data point ($\mathscr{D}_{\mathrm{sh}}=0.141$). We want to remark here, that it took the reinforcement learning agent only 400 episodes to achieve this result compared to the baseline set that is grounded on the generation of 76980 random samples. The reinforcement learning algorithm can therefore be seen as being data efficient.
Nevertheless, in this study, it seems that a lower bound is existing that is difficult to overcome by the MEG-SGGPO approach. In general, this can have two reasons: First, the forming process is unable to produce textures that are closer to the goal texture identified by the SMTLO approach. Yet, the SMTLO approach already addresses this issue by enforcing microstructures to remain within the region delineated by the known microstructures from the training data set via the term $\mathscr{V}_{\mathrm{valid}}\mathscr{C}_{\mathrm{valid}}$ in Equation (12). We expected that a stronger enforcement leads to textures that are better reachable by the process. This, however, can lead to a lower accuracy in terms of obtaining the desired material properties, due to the change in the other weights in the objective function of the SMTLO optimizer. Second, the MEG-SGGPO approach identified a local optimum and got stuck. This can be mitigated by longer MEG-SGGPO runs with optimized hyperparameters. Longer runs, however, have not shown to yield significant improvements in the presented study. As future work, it is desirable to incorporate the knowledge contained in the generated training samples for SMTLO into MEG-SGGPO a priori to enhance its performance.
5 Conclusion
In summary, the machine learning-based approach presented in this study enables accelerated materials development by taking into account details of the manufacturing process. This was particularly challenging as for the considered metal forming process a total of $25^{7}$ different processing paths would be possible; we found that
This shows that the applied approach is highly data efficient and capable of effectively optimizing process-structure-property relations end-to-end in manufacturing processes. Moreover, we emphasize that the approach leverages the non-uniqueness of the materials and process design problems, i.e. a diverse set of microstructures is identified for given desired properties and subsequently, the best reachable microstructure is produced. This is a significant advantage when transferring the approach to real manufacturing systems, where often constraints exist, which may exclude microstructures from being producible.
Nonetheless, the application still needs to be validated in real manufacturing systems. In our specific case, the amount of required training data is relatively high as compared to what is available in typical production workflows. Thus, for application in real-world manufacturing, a future scenario could involve pre-training the machine learning models with a numerical simulation that serves as a digital twin of the process and the material.